Giorgio offers the person who purchases a 250 ml bottle of Allure two free gifts, either an umbrella, a 30 ml bottle of Midnight, a feminine shaving kit, a raincoat or a pair of rain boots. If you purchased Allure what is the probability you selected at random an umbrella and a shaving kit in that order?

Select one:

a. 0.00

b. 1.00

c. 0.05

d. 0.20
Answer #1

Gift choices = 5

We need to choose 2 out of 5 so total number of possibilities = 5C2 = (5*4)/2 = 10

But when order matters, it would be 10*2 = 20

Hence,

P(Required combination) = 1/20 = 0.05

Option C is correct.